OpenStudy (anonymous):

what fraction is the same as 62/8

OpenStudy (anonymous):

Hint: Prime factorizations \(62=31\cdot 2\) and \(8=2^3\).

OpenStudy (anonymous):

ughh which is im so not good at math

OpenStudy (anonymous):

This means: \[\frac{62}{8}=\frac{31 \cdot 2}{2^3}=\frac{31\cdot 2}{2 \cdot 2 \cdot 2}=\cdots\]